OpenStudy (anonymous):

Solve the equation: e^2x - e^x - 20 = 0

OpenStudy (anonymous):

substitute e^x = u and then plug it in the equation and solve it

rvc (rvc):

a quadratic equation \[\rm ax^2+bx+c=0 \\roots~will~be~: x=\large\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

rvc (rvc):

\(\rm Here~e^x=x\)

OpenStudy (anonymous):

but dont you have to use logs to cancel out the e? (or ln)

rvc (rvc):

first find out the root

OpenStudy (anonymous):

and how would I do that?

zepdrix (zepdrix):

You need to first apply your exponent rule: \(\large\rm e^{2x}=(e^x)^2\) That should make it easier to see why a substitution is appropriate.

zepdrix (zepdrix):

\[\large\rm e^{2x} - e^x - 20 = 0\]Becomes,\[\large\rm (e^{x})^2 - e^x - 20 = 0\]And then if you make a substitution,\[\large\rm (\color{orangered}{e^x})^2 - \color{orangered}{e^x} - 20 = 0\]it becomes a simple quadratic, ya?\[\large\rm (\color{orangered}{u})^2 - \color{orangered}{u} - 20 = 0\]

zepdrix (zepdrix):

Just in case you weren't following what the other guys were saying ^

rvc (rvc):

yup so a=1 b=-1 c=20

rvc (rvc):

c= -20*

zepdrix (zepdrix):

I'm not sure if you'll need the fancy quadratic formula for this one :) It looks like it will factor nicely!

rvc (rvc):

yup but this is bit easy to explain lol

OpenStudy (anonymous):

yes I understand changing it to 'u' but where do you go from there? I'm just unsure of how to simplify it?

OpenStudy (anonymous):

@zepdrix @rvc

zepdrix (zepdrix):

\[\large\rm u^2-1u-20=0\]Find factors of -20 which add to -1. \(\rm -10\cdot2=-20\) \(\rm -10+2\ne-1\) Hmm these factors did not work out. \(\rm -20\cdot1=-20\) \(\rm -20+1\ne-1\) Hmm these either...

OpenStudy (anonymous):

itd -5 and 4

zepdrix (zepdrix):

Try other factors of -20. If you're really uncomfortable with factoring, you can use the Quadratic Formula as Rvc suggested,\[\large\rm u=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]Where our quadratic is given by,\[\large\rm au^2+bu+c=0\]So we identify that: \(\rm a=1\) \(\rm b=-1\) \(\rm c=-20\)

zepdrix (zepdrix):

\[\large\rm u^2-1u-20=0\]Ok good. You can do it the shortcut way if you like,\[\large\rm (u-5)(u+4)=0\]Or you understand how we got to that point, you can factor by grouping. Rewrite -1u as the sum of -5u and 4u,\[\large\rm u^2-5u+4u-20=0\]Factor a u out of the first two terms,\[\large\rm u(u-5)+4u-20=0\]Factor a 4 out of the last two terms,\[\large\rm u(u-5)+4(u-5)=0\]Factor a (u-5) out of everything,\[\large\rm (u-5)(u+4)=0\]

zepdrix (zepdrix):

From that point, apply your Zero-Factor Property, setting each individual factor equal to zero, (because a product can only be zero if one of the factors is zero),\[\large\rm u-5=0\qquad\qquad\qquad\qquad\qquad u+4=0\]And from this point, you can undo your substitution,\[\large\rm e^x-5=0\qquad\qquad\qquad\qquad\qquad e^x+4=0\]and apply logs and all that fun stuff to solve for x.

OpenStudy (anonymous):

ok so you would bring the -5 and 4 to the other side and then make it (ln)5 and (ln)-4 right?

zepdrix (zepdrix):

Good good good. Check to make sure that both of your solutions make sense.

zepdrix (zepdrix):

\[\large\rm \ln(5)\approx 1.609\]\[\large\rm \ln(-4)=?\]

OpenStudy (anonymous):

yeah ln(-4) us undefined so the answer is 1.609. thank you!

zepdrix (zepdrix):

np